# Inferring Latent Temporal Sparse Coordination Graph for Multi-Agent Reinforcement Learning

### v1

@@ I Introduction
While these methods balance computational efficiency with effective agent interaction and complex decision-making, in the real world, agents should not only consider their own observations but also take into account the situations of others when taking action [7]. Effective cooperation among agents emerges as a pivotal factor in achieving specific objectives. This cooperation can be assumed to have some latent graph structures [8]. Since the graph is not explicitly given, the inference of meaningful dynamic graph topology has been a persistent challenge.
@@
Figure 2: The framework of LTS-CG. LTS-CG samples a sparse graph from an agent-pair probability matrix generated by agents’ observation trajectories. We further learn the graph to encode two essential characteristics: Predict-Future enables agents to predict future steps effectively, and Infer-Present empowers every partially observed agent to deduce the current state using graph information. The learned temporal graph structure captures past experiences, while current observations determine edge weights. This allows agents to exchange knowledge during policy learning and gain historical and present perspectives for effective cooperation. Graph learning and agent training occur simultaneously in an end-to-end fashion.
@@ II Preliminaries
We focus on cooperative multi-agent tasks modelled as a Decentralized Partially Observable Markov Decision Process (Dec-POMDP) [16] consisting of a tuple $\langle\mathcal{I},\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n},P,\{\mathcal{O}^{i}\}_{i=1}^{n},\{\sigma^{i}\}_{i=1}^{n},R,\gamma\rangle$, where $\mathcal{I}$ is the finite set of $n$ agents, $s\in\mathcal{S}$ is the true state of the environment. At each time step, each agent $i$ observes the state partially by drawing observation $o^{i}_{t}\in\mathcal{O}^{i}$ and selects an action $a^{i}_{t}\in\mathcal{A}^{i}$ according to its own policy $\sigma^{i}$. Individual actions form a joint action $\boldsymbol{a}=(a_{1},...,a_{n})$, which leads to the next state $s^{\prime}$ according to the transition function $P(s^{\prime}|s,\boldsymbol{a})$ and a reward $R(s,\boldsymbol{a})$ shared by all agents. Each agent has local action-observation history $\tau_{i,t}=(o_{i,0},a_{i,0},...,o_{i,t-1},a_{i,t-1},o_{i,t})\in(\mathcal{O}^{i}\times\mathcal{A}^{i})^{t}\times\mathcal{O}^{i}$. This paper considers episodic tasks yielding episodes $(s_{0},\{o^{i}_{0}\}_{i=1}^{n},\boldsymbol{a}_{0},r_{0},...,s_{T},\{o^{i}_{T}\}_{i=1}^{n})$ of varying finite length $T$. Agents learn to collectively maximize the global return $Q_{tot}(s,\boldsymbol{a})=\mathbb{E}_{s_{0:T},a_{0:T}}\left[\sum_{t=0}^{T}\gamma^{t}R\left(s_{t},\boldsymbol{a}_{t}\right)\mid s_{0}=s,\boldsymbol{a}_{0}=\boldsymbol{a}\right]$, where $\gamma\in[0,1)$ is the discount factor.
@@ III-A Graph-based MARL
As the graph itself is not explicitly given, inferring graph topology remains a critical prerequisite for training MARL. From the perspective of graph structure, existing methods for graph inference can be broadly categorized into three types: (a) creating fully connected unweighted graphs by directly linking all nodes/agents explicitly, such as DGN [24], PIC [9] and DCG [10] or implicitly such as MAAC [25], ROMA [26]; (b) employing attention mechanisms to calculate fully connected weighted graphs, such as GraphMIX [11] and DICG [12]; (c) designing drop-edge criteria to generate sparse weighted graphs, such as random drop edges in G2ANet [21], select sparse graph from candidate set in SOP-CG [13] and drop edges based on variance of payoff functions in CASEC [14].
Despite this progress, these methods exhibit the following limitations: one is that they primarily focus on one-step observations and fail to consider the value of historical trajectory data, which more accurately represents agents’ behaviours and is more meaningful to help to learn policies [15]; another is that the computation-intensive nature of action-pair calculations in coordination graphs (CG) [8] poses significant scalability challenges, which becomes increasingly problematic as the number of agents and actions increases (See: V-A).
@@ IV-A1 Sparse graph construction
$\min_{w}\quad\mathrm{E}_{A\sim\operatorname{Ber}(\theta(w))}\left[\mathcal{L}\left(A,w,\mathcal{O}_{T}\right)\right].$ (3)
Here, $\mathcal{O}_{T}=\{{\mathcal{O}^{i}_{T}\}_{i=1}^{n}}$, and $\mathcal{O}^{i}_{T}=\{o^{i}_{0},...,o^{i}_{T}\}$ denotes the observation trajectory for agent $i$ over the time steps $T$. Each element of $A$ is sampled from a Bernoulli distribution $\operatorname{Ber}(\theta(w))$, with $\omega$ denoting the trainable weight. In Eq. (3), the adjacent probability $\theta$ is absorbed together with the GNNs parameters $w$, making the gradient computation more efficient and having better scalability [32]. In the following, we give the details about how to infer the inter-agent sparse graph $A$ and how to define the graph learning loss function $\mathcal{L}$.
@@
where $\text{FC}(\cdot)$ is a fully connected layer and $\text{CONV}(\cdot)$ is the convolution layer performed along the temporal dimension. This convolutional layer plays a crucial role in capturing each agent’s latent behaviour patterns over time, enhancing the model’s ability to discern dynamic and temporal patterns in the agents’ interactions. Then the agent-pair predictor $f_{ap}(\cdot)$ utilize the temporal dependencies of every agent-pair ($z^{i}$ and $z^{j}$) to calculate adjacent probability $\theta_{ij}$ as follows
@@ IV-A2 Meaningful graph learning
Motivated by the idea that the graph should enhance the agents’ understanding of other agents and the environment, we further learn the graph to encode the following two essential characteristics.
@@
Figure 3: Performance of our method and baselines on six maps of the StarCraft II benchmark [40]. Y-axis is the test winning rate of the game. X-axis is training steps
@@ IV-B Cooperative MARL with LTS-CG
In our design, LTS-CG facilitates simultaneous graph inference and multi-agent policy learning for efficient end-to-end training. Initially, with the buffer’s setup, we store the inter-agent graph with full connectivity. This allows agents to start the game with the capability to cooperate and make decisions. As training progresses, LTS-CG dynamically learns and updates the graph’s structure within the buffer. During testing, the graph, retrieved from the buffer, is updated with edge weights based on current observations. This temporal sparse graph is then utilized for message calculation and dissemination, ensuring agents have access to the latest information. Such an approach guarantees effective decision-making and cooperation throughout the training process.
@@
$\underset{\mathbf{a}}{\operatorname{argmax}}Q_{tot}(\boldsymbol{\tau},\mathbf{a})=\left(\begin{array}[]{c}\operatorname{argmax}_{a_{1}}Q_{1}\left(\tau_{1},a_{1},m_{1}\right)\\ \vdots\\ \operatorname{argmax}_{a_{n}}Q_{n}\left(\tau_{n},a_{n},m_{n}\right)\end{array}\right).$ (15)
@@ V Experiments
In this section, we design experiments to answer the following questions: (1) How does LTS-CG compare in performance with state-of-the-art methods on complex cooperative multi-agent tasks? (See: V-A) (2) Is the utilization of trajectories for learning the coordination graph more effective than relying on one-step observations? (See: V-B1) (3) Does having the Predict-Future and Infer-Present characteristics improve the performance of LTS-CG? (See: V-B2) (4) What are the effects of varying the weights for $\mathcal{L}_{g}$ on the experimental outcomes? (See: V-B3)
To answer the above questions, we conduct our experiments on StarCraft II benchmark [40], which consists of different maps with varying numbers of agents. The experiments included scenarios with a minimum of eight agents, comprising both homogeneous and heterogeneous agent setups. All the experiments are carried out with difficulty=7 and repeated with 5 random seeds. We employ distinct 2-layer GNNs as specified in Eq.(11) to facilitate the acquisition of the Infer-present characteristic and to compute the knowledge exchanged during agents’ cooperation. The graph loss $\lambda$, the character-balance weight $b$ and $c$ in Eq (16) are set to 1. The hyperparameters used for MARL in StarCraft II are given in Tab. II. Experiences are stored in a first-in-first-out (FIFO) replay buffer and during the training phase. The experiments are finished with Intel(R) Xeon(R) E-2288G CPU and Tesla V100-PCIE-32GB GPU. The software that we use for experiments is Python 3.7.13, PyTorch 1.13.1, PyYAML 6.0, numpy 1.21.5 and CUDA 11.6.
Maps
Agent
Num.
Enemy
Num.
Action
Num.
Episode
Limit
8m 8 8 14 120
25m 25 25 31 150
3s5z 8 8 14 150
1c3s5z 9 9 15 180
8m_vs_9m 8 9 15 120
10m_vs_11m 10 11 17 150
27m_vs_30m 27 30 36 180
MMM2 10 12 18 180
TABLE I: Detailed information of each map we used in StarCraft II benchmark.
Hyperparameter Value
Batch-size 32
Replay memory size 5000
discount factor $\gamma$ 0.99
Optimizer RMSProp
Learning rate $5\times 10^{-4}$
optim_alpha 0.99
optim_eps $1\times 10^{-5}$
Gradient-norm-clip 10
Action-selector $\epsilon$-greedy
$\epsilon$-start 1.0
$\epsilon$-finish 0.05
$\epsilon$-anneal-time 50000 steps
target update interval 200
TABLE II: The hyperparameters used in StarCraft II.
Method Graph type Edge Data in learning graph
QMIX $\times$ $\times$ $\times$
DCG Complete Unweighted One-step
DICG Complete Weighted One-step
SOP-CG Sparse Unweighted One-step
CASEC Sparse Weighted One-step
LTS-CG Sparse Weighted Trajectories
TABLE III: Comparison of different experiment methods in terms of graph type, edge representation, and data used for learning graph.
Figure 4: Performance comparison on the 25m and 27m_vs_30m maps. Due to the high computational complexity, SOP-CG and CASEC could not complete the experiments on both maps. DCG could not finish the experiment on the later map.
@@ V-A Performance comparison on StarCraft II
V-A1 Details for comprised methods
We utilize several state-of-the-art baseline algorithms for our experiments. Each method’s graph type, edge representation, and group utilization are summarised in Tab. III. Below, we provide a brief introduction of each method and the detailed settings we used:
@@ V-A1 Details for comprised methods
V-A2 Results
Fig. 3 presents the results of our method compared to the performance of other algorithms on six different maps. The experimental results clearly demonstrate the superiority of our approach LTS-CG across all scenarios (shown in orange). Firstly, our method exhibited faster convergence than the compared methods on all six maps in the early stages of training (below 0.6 mil for 8m, 2 mil for MMM2, and 1 mil for other maps). This indicates that our approach enables the agents to quickly learn effective cooperative strategies and achieve high-performance levels. Moreover, our method demonstrated a smaller standard deviation in performance compared to the other methods, such as CASEC in 3s5z, DICG in 8m_vs_9m and DCG in 10m_vs_11m. The reduced variability suggests that our approach consistently produces reliable and stable cooperative behaviours, resulting in more predictable and robust performance across different maps. Notably, our method achieved consistent and competitive performance across all six maps. This indicates that our approach generalizes well and is capable of adapting to various environmental conditions and agent configurations. The ability to achieve good results consistently is essential for real-world applications of multi-agent systems.
@@ V-A2 Results
Figure 5: Performance comparison on the 3s5z and 10m_vs_11m. OneStepObs-c and OneStepObs-s utilize a one-step observation to generate a fully connected graph and a sparse graph separately. LTS-CG$(w/o\mathcal{L}_{g})$ investigates the effectiveness of using trajectories by omitting the Predict-Future and Infer-Present characteristics.
Figure 6: Evaluate the effectiveness of the different latent temporal sparse graph learning strategies on two maps. LTS-CG$(w/o\mathcal{L}g)$ excludes both Predict-Future and Infer-Present characteristics. LTS-CG$(\mathcal{L}{pre})$ and LTS-CG$(\mathcal{L}_{inf})$ only incorporate the Predict-Future or Infer-Present characteristic into the learning process separately.
@@
V-B Ablation study
V-B1 Trajectory Graph Learning vs One-Step Observations

### v2

@@ I Introduction
While these methods balance computational efficiency with effective agent interaction and complex decision-making, in the real world, agents should not only consider their own observations but also take into account the situations of others when taking action [7]. Effective cooperation among agents emerges as a pivotal factor in achieving specific objectives. This cooperation can be assumed to have some latent graph structures [8]. Since the agent graph is not explicitly given, the inference of meaningful dynamic graph topology has been a persistent challenge.
@@
Figure 2: The framework of LTS-CG. LTS-CG consists of two key modules: Inter-Agent Sparse Graph Learning and Cooperative MARL. The former follows an encoder-decoder framework: the encoder generates the sparse graph structure, while the decoder—guided by two graph loss functions—learns Predict-Future for anticipating future steps and Infer-Present for deducing current states. The temporal graph structure integrates past experiences and adjusts edge weights based on current observations. This graph is then fed into the attention-based graph convolution of the Cooperative MARL module, enabling knowledge exchange for effective coordination. Graph learning and agent training occur end-to-end.
@@ II Preliminaries
We focus on cooperative multi-agent tasks modelled as a Partially Observable Markov Decision Process (POMDP) [16] consisting of a tuple $\langle\mathcal{I},\mathcal{S},\{\mathcal{A}^{i}\}_{i=1}^{n},P,\{\mathcal{O}^{i}\}_{i=1}^{n},\{\sigma^{i}\}_{i=1}^{n},R,\gamma\rangle$, where $\mathcal{I}$ is the finite set of $n$ agents, $s\in\mathcal{S}$ is the true state of the environment. At each time step, each agent $i$ observes the state partially by drawing observation $o^{i}_{t}\in\mathcal{O}^{i}$ and selects an action $a^{i}_{t}\in\mathcal{A}^{i}$ according to its own policy $\sigma^{i}$. Individual actions form a joint action $\boldsymbol{a}=(a_{1},...,a_{n})$, which leads to the next state $s^{\prime}$ according to the transition function $P(s^{\prime}|s,\boldsymbol{a})$ and a reward $R(s,\boldsymbol{a})$ shared by all agents. Each agent has local action-observation history $\tau_{i,t}=(o_{i,0},a_{i,0},...,o_{i,t-1},a_{i,t-1},o_{i,t})\in(\mathcal{O}^{i}\times\mathcal{A}^{i})^{t}\times\mathcal{O}^{i}$. This paper considers episodic tasks yielding episodes $(s_{0},\{o^{i}_{0}\}_{i=1}^{n},\boldsymbol{a}_{0},r_{0},...,s_{T},\{o^{i}_{T}\}_{i=1}^{n})$ of varying finite length $T$. Agents learn to collectively maximize the global return $Q_{tot}(s,\boldsymbol{a})=\mathbb{E}_{s_{0:T},a_{0:T}}\left[\sum_{t=0}^{T}\gamma^{t}R\left(s_{t},\boldsymbol{a}_{t}\right)\mid s_{0}=s,\boldsymbol{a}_{0}=\boldsymbol{a}\right]$, where $\gamma\in[0,1)$ is the discount factor.
@@ III-A Graph-based MARL
As the graph itself is not explicitly given, inferring graph topology remains a critical prerequisite for training MARL. From the perspective of graph structure, existing methods for graph inference can be broadly categorized into three types: (a) creating fully connected unweighted graphs by directly linking all nodes/agents explicitly, such as DGN [24], PIC [9] and DCG [10], or implicitly such as MAAC [25], ROMA [26]; (b) employing attention mechanisms to calculate fully connected weighted graphs, such as GraphMIX [11] and DICG [12]; (c) designing drop-edge criteria to generate sparse weighted graphs, such as random drop edges in G2ANet [21], select sparse graph from candidate set in SOP-CG [13], drop edges based on variance of payoff functions in CASEC [14], and and generate event graphs with certain rules in CAAC [27].
Despite this progress, these methods exhibit the following limitations: one is that they primarily focus on one-step observations and fail to consider the value of historical trajectory data, which more accurately represents agents’ behaviours and is more meaningful to help to learn policies [15]; another is that the computation-intensive nature of action-pair calculations in coordination graphs (CG) [8] poses significant scalability challenges, which becomes increasingly problematic as the number of agents and actions increases (See: V-A1 and V-E).
@@ IV-A1 Sparse Graph Construction
${\color[rgb]{0,0,0}\min_{\omega}\quad\mathrm{E}_{A\sim\operatorname{Ber}(\theta(\omega))}\left[\mathcal{L}\left(A,\omega,\mathcal{O}_{T}\right)\right].}$ (3)
Here, $\mathcal{O}_{T}=\{{\mathcal{O}^{i}_{T}\}_{i=1}^{n}}$, and $\mathcal{O}^{i}_{T}=\{o^{i}_{0},...,o^{i}_{T}\}$ denotes the observation trajectory for agent $i$ over the time steps $T$. Each element of $A$ is sampled from a Bernoulli distribution $\operatorname{Ber}(\theta(\omega))$, with $\omega$ denoting the trainable weight. In Eq. (3), the adjacent probability $\theta$ is absorbed together with the GNNs parameters $\omega$, making the gradient computation more efficient and having better scalability [33]. In the following, we give the details about how to infer the inter-agent sparse graph $A$ and how to define the graph learning loss function $\mathcal{L}$.
@@
where $\text{FC}(\cdot)$ is a fully connected layer and $\text{CONV}(\cdot)$ is the convolution layer performed along the temporal dimension. Since the episodes may end before reaching the maximum time step, zeros are padded in $\mathcal{O}^{i}_{T}$ to ensure consistent input length across episodes. This convolutional layer plays a crucial role in capturing each agent’s latent behaviour patterns over time, enhancing the model’s ability to discern dynamic and temporal patterns in the agents’ interactions. Then the agent-pair predictor $f_{ap}(\cdot)$ utilize the temporal dependencies of every agent-pair ($z^{i}$ and $z^{j}$) to calculate adjacent probability $\theta_{ij}$ as follows
@@ IV-A2 Meaningful Graph Learning
Motivated by the idea that the graph should enhance the agents’ understanding of other agents and the environment, we further learn the graph to have the following two essential characteristics.
@@ IV-B Cooperative MARL with LTS-CG
In our LTS-CG design, graph inference and multi-agent policy learning are integrated for efficient end-to-end training. The Inter-Agent Sparse Graph Learning module follows an encoder-decoder framework: the encoder generates the sparse graph structure, while the decoder—guided by the two graph loss functions mentioned earlier—refines the encoder’s weights.
At the start of training, the buffer stores a fully connected inter-agent graph, allowing agents to cooperate and make informed decisions from the outset. As training progresses, our method learns a temporally sparse graph, which is stored in the buffer and reused in subsequent training iterations, thereby accelerating the learning process. During testing, only the encoder is used to generate the graph structure. The resulting graph is then fed into an attention-based graph convolution part of Cooperative MARL module, dynamically adjusting edge weights at each time step. This temporal sparse graph ensures that agents always have access to the most up-to-date information for effective decision-making and coordination throughout the training process.
@@
$\underset{\mathbf{a}}{\operatorname{argmax}}\quad Q_{tot}(\boldsymbol{\tau},\mathbf{a})=\left(\begin{array}[]{c}\operatorname{argmax}_{a_{1}}Q_{1}\left(\tau_{1},a_{1},m_{1}\right)\\ \vdots\\ \operatorname{argmax}_{a_{n}}Q_{n}\left(\tau_{n},a_{n},m_{n}\right)\end{array}\right).$ (15)
@@
Figure 3: Performance of our method and baselines on six maps of the StarCraft II benchmark [41]. The Y-axis is the test winning rate of the game. The X-axis is the training steps.
@@ V Experiments
In this section, we design experiments to answer the following questions: (1) How does LTS-CG compare in performance with graph-based methods on complex cooperative multi-agent tasks? (See: V-A1) (2) How does LTS-CG compare in performance with non-graph methods? (See: V-A2) (3) How does LTS-CG perform across a variety of scenarios? (See: V-B) (4) Is the utilization of trajectories for learning the coordination graph more effective than relying on one-step observations? (See: V-C1) (5) Is sampling from the Attention Matrix necessary? (See: V-C2) (6) Does having the Predict-Future and Infer-Present characteristics improve the performance of LTS-CG? (See: V-C3) (7) What are the effects of varying the weights for $\mathcal{L}_{g}$ on the experimental outcomes? (See: V-C4)
To answer the above questions, our experiments involve the following three environments:
  • StarCraft II benchmark (SMAC) [41]: consists of different maps with varying numbers of agents. Our experiments included scenarios with a minimum of eight agents, comprising both homogeneous and heterogeneous agent setups. All the experiments are carried out with difficulty=7.
  • Tag (MPE) [18]: is a task based on the particle world environment. In this scenario, a group of agents chases several adversaries on a map containing three randomly generated obstacles. The agents receive a global reward for each collision with an adversary. The adversaries move faster, making it crucial for the agents to collaborate effectively to surround them. We tested the common setup of 10 agents chasing 3 adversaries, and we further extended this to 20 agents chasing 5 adversaries to evaluate the scalability of different methods.
  • Gather: is an extension of the Climb Game [42]. In the original Climb Game, each agent has three possible actions, $A=\{a_{0},a_{1},a_{2}\}$. Action $a_{0}$ yields no reward unless all agents choose it, at which point it provides a high reward. The other two actions are sub-optimal but can yield positive rewards without requiring perfect coordination. We followed the setup from the MULTI-AGENT COORDINATION BENCHMARK (MACO) [14] for our experiments.
We employ distinct 2-layer GNNs as specified in Eq.(11) to facilitate the acquisition of the Infer-present characteristic and to compute the knowledge exchanged during agents’ cooperation. The graph loss $\lambda$, the character-balance weight $b$ and $c$ in Eq (16) are set to 1. Experiences are stored in a first-in-first-out (FIFO) replay buffer during the training phase, and all settings are repeated with five random seeds for consistency. The experiments are finished with Intel Xeon Gold 6226R CPUs and NVIDIA Quadro RTX 8000 GPUs (48 GB) GPU. The software that we use for experiments is Python 3.7.13, PyTorch 1.13.1, PyYAML 6.0, numpy 1.21.5 and CUDA 11.6. More experimental details and our implementation can be found at https://github.com/Wei9711/LTSCG
@@ V-A Performance Comparison on StarCraft II
V-A1 Comparison with Graph-based Methods
We utilize several state-of-the-art baseline algorithms for our experiments. Below, we provide a brief introduction of each method and the detailed settings we used:
@@ V-A1 Comparison with Graph-based Methods
Figure 4: Performance comparison on the 25m and 27m_vs_30m maps. Due to the high computational complexity, SOP-CG and DCG could not complete 2 million steps within a week, and CASEC exceeded the 48 GB GPU memory limit.
Results: Fig. 3 presents the results of our method compared to the performance of other algorithms on six different maps. The experimental results clearly demonstrate the superiority of our approach LTS-CG across all scenarios (shown in black). Firstly, our method exhibited faster convergence than the compared methods on all six maps in the early stages of training (below 0.6 mil for 8m, 2 mil for MMM2, and 1 mil for other maps). This indicates that our approach enables the agents to quickly learn effective cooperative strategies and achieve high-performance levels. Moreover, our method demonstrated a smaller standard deviation in performance compared to the other methods, such as CASEC in 3s5z, DICG in 8m_vs_9m and DCG in 10m_vs_11m. The reduced variability suggests that our approach consistently produces reliable and stable cooperative behaviours, resulting in more predictable and robust performance across different maps. Notably, our method achieved consistent and competitive performance across all six maps. This indicates that our approach generalizes well and is capable of adapting to various environmental conditions and agent configurations. The ability to achieve good results consistently is essential for real-world applications of multi-agent systems.
@@
Figure 5: Performance comparison of non-graph-based methods on 3s5z and 8m_vs_9m.
@@
Figure 6: Performance comparison of different methods on the TAG and Gather scenarios. The TAG scenario involves agents chasing adversaries on a map with obstacles, showing results for 10 agents and 3 adversaries and 20 agents and 5 adversaries The Gather scenario is an extension of the Climb Game where precise coordination yields higher rewards.
Figure 7: Performance comparison on two maps to evaluate whether utilizing trajectories is more effective than relying solely on one-step observations.
V-A2 Comparison with Non-graph-based Methods
In this subsection, we further include several non-graph-based methods for comparison with our proposed method, as these are widely accepted as benchmarks and are frequently used to assess the performance of new algorithms in the SMAC environment:
  • DDN  [^6] https://github.com/j3soon/dfac[43] uses distributional reinforcement learning to factorize the value function by modelling utility functions as random variables and applying a quantile mixture. We standardized the agent dimensions to 64 instead of the original 256 for consistency across methods.
  • MAPPO  [^7] https://github.com/marlbenchmark/on-policy [44] is a widely used method for cooperative multi-agent reinforcement learning, which has been shown to perform well in both continuous and discrete action spaces.
  • MADDPG  [^8] https://github.com/uoe-agents/epymarl[18] is designed for mixed cooperative-competitive environments and leverages centralized training with decentralized execution.
[…]
Figure 8: Performance comparison on two maps to evaluate whether sampling the graph from the attention matrix is more effective than not sampling.
V-B Performance Comparison on Tag and Gather
We further evaluated the performance of different methods on the TAG and Gather scenarios. Fig. 6 shows the results for 10 agents and 3 adversaries and the results for 20 agents and 5 adversaries. The Gather scenario is an extended version of the Climb Game, where precise coordination is essential for achieving higher rewards.
Our proposed method, LTS-CG, consistently demonstrates competitive performance across both scenarios. In the TAG scenario, LTS-CG scales efficiently from 10 to 20 agents, maintaining robust performance. Notably, when the number of agents increases to 20, two state-of-the-art methods, SOP-CG and CASEC, fail to complete training within 7 days under the same experimental conditions. The performance of DCG drops significantly when scaling from 10 to 20 agents. In the Gather scenario, LTS-CG exhibits the best test return after 0.8 million steps and demonstrates a promising convergence speed compared to DICG. These results illustrate the scalability, robustness, and efficiency of LTS-CG across a variety of environments and agent numbers, proving its capability to handle complex multi-agent scenarios.
Figure 9: Evaluate the effectiveness of the different latent temporal sparse graph learning strategies on SMAC and Gather.
V-C Ablation Study
V-C1 Trajectory Graph Learning vs One-Step Observations